# Arithmetic dynamics and Generalized Fermat’s conjecture

### v2

@@
Arithmetic dynamics and Generalized Fermat’s conjecture
Atsushi MORIWAKI School of General Education, Chubu University, 1200 Matsumoto-cho, Kasugai-shi, Aichi 487-8501, JAPAN moriwaki@fsc.chubu.ac.jp
(Date: 14/Jan/2025)
@@ 1. Introduction
Let $K$ be an arithmetic function field (i.e., $K$ is finitely generated over $\mathbb{Q}$}) and $S=((\Omega,\mathcal{A},\nu),\{|\raisebox{1.72218pt}{.}|_{\omega}\}_{\omega\in\Omega})$ be a proper adelic structure of $K$ such that $S$ has Northcott’s property and $\nu(\mathcal{A})\not\subseteq\{0,1\}$ (see [3, Proposition 2.7.8] for the existence of such adelic structure). Let $X$ be a geometrically integral projective scheme over $K$ and $L$ be an ample line bundle on $X$. Fix a positive integer $N_{0}$. Let $F=\{f_{N}\}_{N=N_{0}}^{\infty}$ be a sequence of endomorphisms of $X$ with the following properties:
@@
The sequence $F$ is called a system of endomorphisms compatible with $L$. We say $F$ is multiplicative (resp. additive) if $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{NN^{\prime}}$ (resp. $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{N+N^{\prime}}$) for all $N,N^{\prime}\geqslant N_{0}$. For example, the cases of Example 3.2 and Example 3.3 are multiplicative, and the case of Example 3.4 is additive. We can assign the height function $h_{F}$ on $X(K)$ to $F$ such that
@@
If there is $N_{1}\geqslant N_{0}$ such that $Y_{N}(K)$ is finite for any integer $N\geqslant N_{1}$, then does there exist $N_{2}\geqslant N_{0}$ such that $Y_{N}$ has Fermat’s property for any $N\geqslant N_{2}$?
@@
    If $F$ is multiplicative and there is $p_{0}$ such that $Y_{p}(K)$ is finite for any prime $p$ with $p\geqslant p_{0}$, then
    $\lim_{m\to\infty}\frac{\{N_{0}\leqslant N\leqslant m\mid\text{$F_{N}$ has Fermat's property over $K$}\}}{m}=1.$
@@ 2. Endomorphism compatible with an ample line bundle
Let $f:X\to X$ be an endomorphism over $K$. Let $L$ be an ample line bundle on $X$. We say $f$ is compatible with $L$ if $\alpha:f^{*}(L)\simeq L^{\otimes d(f)}$ for some positive integer $d(f)$.
@@
From now on, we assume that $d(f)$ is greater than or equal to 22.
@@
(5) Fix isomorphisms $\alpha:f^{*}(L)\simeq L^{\otimes d(f)}$ and $\beta:g^{*}(L)\simeq L^{\otimes d(g)}$. Then, by [3, Proposition 9.3.4], we can find $r\in K^{\times}$ such that $|\raisebox{1.72218pt}{.}|_{\varphi_{g,\beta,\omega}}=|r|_{\omega}^{-1/(d(f)-1)(d(g)-1)}|\raisebox{1.72218pt}{.}|_{\varphi_{f,\alpha,\omega}}$. Thus the first assertion follows. Moreover,
$(f\circ g)^{*}(L)=f^{*}(g^{*}(L))\simeq f^{*}(L^{\otimes d(g)})=f^{*}(L)^{\otimes d(g)}\simeq(L^{\otimes d(f)})^{\otimes d(g)}=L^{\otimes d(f)d(g)}.$
Thus the second assertion follows.
@@ 3. Generalized Fermat’s conjecture
    The height function $h_{f_{N}}$ ($N\geqslant N_{0})$ does not depend on NN, so they are denoted by $h_{F}$.
@@
In Example 3.2, we assume that $K={\mathbb{Q}}$, $n=2$ and $Y$ is defined by $X_{0}+X_{1}-X_{2}=0$. Note that $Y_{N}$ is given by $X_{0}^{N}+X_{1}^{N}-X_{2}^{N}$. If $Y_{N}$ has Fermat’s property, then, for $(x_{0}:x_{1}:x_{n})\in Y_{N}(K)$, we can find $t\in{\mathbb{Q}}^{\times}$ such that $tx_{0},tx_{1},tx_{2}\in\{0,1,-1\}$. Thus we can easily see that $x_{0}x_{1}x_{2}=0$, and hence Fermat’s conjecture for NN follows.
@@
Proof of Theorem 1.2.
$d_{m}a(h_{F})>\max\{h_{F}(y)\mid y\in S\}$
for all $m\in{\mathbb{Z}}_{\geqslant m_{0}}$.
(3) For any prime number $p\geqslant p_{0}$, we set $S=Y_{p}(K)$. If $x\in Y_{mp}(K)$, then one has $f_{m}(x)\in Y_{p}(K)$ because $f_{p}(f_{m}(x))=f_{mp}(x)\in Y(K)$. Moreover, $f_{m}\circ f_{N_{0}}=f_{N_{0}}\circ f_{m}$. Thus, by Lemma 2.2, $Y_{mp}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}$ for $m\in{\mathbb{Z}}_{\geqslant m_{0}}$. Therefore, by [8, Lemma 5.16], we have the assertion. ∎

### v3

@@
Abstract.
In this short paper, we would like to propose generalized Fermat’s conjecture over an arithmetic function field in the framework of arithmetic dynamics. We also give several evidences for the conjecture.
I would like to thank Prof. Kawaguchi for his helpful comments.
@@ 1. Introduction
Let $K$ be an arithmetic function field (i.e., $K$ is finitely generated over $\mathbb{Q}$}) and $S=((\Omega,\mathcal{A},\nu),\{|\raisebox{1.72218pt}{.}|_{\omega}\}_{\omega\in\Omega})$ be a proper adelic structure of $K$ such that $S$ has Northcott’s property and $\nu(\mathcal{A})\not\subseteq\{0,1\}$ (for the definition of “adelic structure” and its related terminology, see [1]). Let $X$ be a geometrically integral projective scheme over $K$ and $L$ be an ample line bundle on $X$. Fix a positive integer $N_{0}$. Let $F=\{f_{N}\}_{N=N_{0}}^{\infty}$ be a sequence of endomorphisms of $X$ with the following properties:
@@
The sequence $F$ is called a system of endomorphisms compatible with $L$. We say $F$ is multiplicative (resp. additive) if $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{NN^{\prime}}$ (resp. $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{N+N^{\prime}}$) for all $N,N^{\prime}\geqslant N_{0}$. For example, the cases of Example 3.2, Example 3.3, Example 3.4 and Example 3.5 are multiplicative, and the case of Example 3.6 is additive. We can assign the height function $h_{F}$ on $X(K)$ to $F$ such that
@@
If there is $N_{1}\geqslant N_{0}$ such that $Y_{N}(K)$ is finite for any integer $N\geqslant N_{1}$, then does there exist $N_{2}\geqslant N_{0}$ such that $Y_{N}$ has Fermat’s property over $K$ for any $N\geqslant N_{2}$? For a multi-indexed version, see Conjecture 4.4.
@@
    If $F$ is multiplicative and there is a positive integer $p_{0}$ such that $Y_{p}(K)$ is finite for any prime $p$ with $p\geqslant p_{0}$, then
    $\lim_{m\to\infty}\frac{\#\{N_{0}\leqslant N\leqslant m\mid\text{$F_{N}$ has Fermat's property over $K$}\}}{m}=1.$
@@ 2. Endomorphism compatible with an ample line bundle
Let $f:X\to X$ be an endomorphism over $K$. Let $L$ be an ample line bundle on $X$. We say $f$ is compatible with $L$ if there is an isomorphism $\alpha:f^{*}(L)\simeq L^{\otimes d(f)}$ for some positive integer $d(f)$.
@@
From now on, we assume that $d(f)$ is greater than or equal to $2$.
@@
(5) Fix isomorphisms $\alpha:f^{*}(L)\simeq L^{\otimes d(f)}$ and $\beta:g^{*}(L)\simeq L^{\otimes d(g)}$. Let us consider the following homomorphisms:
$\begin{cases}\begin{CD}g^{*}(f^{*}(L))@>{\sim}>{{g^{*}(\alpha)}}>g^{*}(L^{\otimes d(f)})@>{\sim}>{\beta^{\otimes d(f)}}>L^{\otimes d(f)d(g)}\end{CD},\\[8.61108pt] \begin{CD}g^{*}(f^{*}(L))=f^{*}(g^{*}(L))@>{\sim}>{f^{*}(\beta)}>f^{*}(L^{\otimes d(g)})@>{\sim}>{\alpha^{\otimes d(g)}}>L^{\otimes d(f)d(g)}\end{CD}.\end{cases}$
Thus we can find $r\in K^{\times}$ such that $\beta^{\otimes d(f)}\circ g^{*}(\alpha)=r\cdot\alpha^{\otimes d(g)}\circ f^{*}(\beta)$. Then, by [3, Proposition 9.3.4], one has $|\raisebox{1.72218pt}{.}|_{\varphi_{g,\beta,\omega}}=|r|_{\omega}^{-1/(d(f)-1)(d(g)-1)}|\raisebox{1.72218pt}{.}|_{\varphi_{f,\alpha,\omega}}$. Thus the first assertion follows. Moreover, by using the above isomorphisms, the second assertion is obtained.
@@ 3. Generalized Fermat’s conjecture
    The height function $h_{f_{N}}$ ($N\geqslant N_{0})$ does not depend on $N$, so they are denoted by $h_{F}$.
@@
Example 3.4.
Here we consider a family of Lattès maps. Let $E$ be an elliptic curve over $K$ and $\pi:E\to{\mathbb{P}}^{1}$ be the morphism such that $\deg(\pi)=2$ and $\pi(P)=\pi(-P)$, that is, ${\mathbb{P}}^{1}=E/\{\pm 1\}$. For an integer $N\geqslant 2$, let $f_{N}:{\mathbb{P}}^{1}\to{\mathbb{P}}^{1}$ be the morphism such that the following diagram is commutative:
[…]
Then $f_{N}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))={\mathcal{O}}_{{\mathbb{P}}^{1}}(N^{2})$ and $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{NN^{\prime}}$. Note that $d_{N}=N^{2}$ and $\{f_{N}\}_{N=2}^{\infty}$ is multiplicative.
Example 3.5.
Let $T_{N}(z)$ be the Chebyshev polynomial of degee $N$, which is given by the recursive relation $T_{N}(z)=zT_{N-1}(z)-T_{N-2}(z)$ ($T_{0}=2$, $T_{1}=z$ and $T_{2}=z^{2}-2$). It can be characterized by $T_{N}(z+z^{-1})=z^{N}+z^{-N}$, so we can see that $T_{N}(T_{N^{\prime}}(z))=T_{N^{\prime}}(T_{N}(z))=T_{NN^{\prime}}(z)$. Thus $\{T_{N}\}_{N=2}^{\infty}$ yields a multiplicative system of endomorphisms. Note that $(1/2)T_{N}(2z)$ is the classical Chebyshev polynomial of degree $N$. For details, see [9, Section 6.2].
@@
In Example 3.2, we assume that $K={\mathbb{Q}}$, $n=2$ and $Y$ is defined by $X_{0}+X_{1}-X_{2}=0$. Note that $Y_{N}$ is given by $X_{0}^{N}+X_{1}^{N}-X_{2}^{N}$. If $Y_{N}$ has Fermat’s property, then, for $(x_{0}:x_{1}:x_{n})\in Y_{N}(K)$, we can find $t\in{\mathbb{Q}}^{\times}$ such that $tx_{0},tx_{1},tx_{2}\in\{0,1,-1\}$. Thus we can easily see that $x_{0}x_{1}x_{2}=0$, and hence Fermat’s conjecture for $N$ follows.
@@
Remark 3.9.
If $\{f_{N}\}_{N=N_{0}}^{\infty}$ is a multiplicative system of endomorphisms compatible with $L$, then, for a fixed integer $m\geqslant 2$, $\{g_{N}\}_{N=N_{1}}^{\infty}$ given by $g_{N}=f_{m^{N}}$ is additive, where $N_{1}$ is an integer with $m^{N_{1}}\geqslant N_{0}$.
@@
Proof of Theorem 1.2.
$d_{m}a(h_{F})>\max\{h_{F}(y)\mid y\in S\}$
for all $m\in{\mathbb{Z}}_{\geqslant m_{0}}$.
(3) For any prime number $p\geqslant p_{0}$, we set $S=Y_{p}(K)$. If $x\in Y_{mp}(K)$, then one has $f_{m}(x)\in Y_{p}(K)$ because $f_{p}(f_{m}(x))=f_{mp}(x)\in Y(K)$. Moreover, $f_{m}\circ f_{N_{0}}=f_{N_{0}}\circ f_{m}$. Thus, by Lemma 4.2, $Y_{mp}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}$ for $m\in{\mathbb{Z}}_{\geqslant m_{0}}$. Therefore, by [8, Lemma 5.16] or Lemma 4.6, we have the assertion. ∎
4. Product case (multi-indexed version)
Let $X_{1},\ldots,X_{n}$ be geometrically integral projective schemes over $K$ and $L_{1},\ldots,L_{n}$ be ample line bundles on $X_{1},\ldots,X_{n}$, respectively. Let $f_{1},\ldots,f_{n}$ be endomorphisms of $X_{1},\ldots,X_{n}$, respectively. We assume that, for each $i\in\{1,\ldots,n\}$, $f_{i}^{*}(L_{i})\simeq L_{i}^{\otimes d_{i}}$ for some integer $d_{i}\geqslant 2$. Let $\varphi_{i}$ be the global compatification of $f_{i}^{*}(L_{i})\simeq L_{i}^{d_{i}}$ for each $i$. Let $X=X_{1}\times\cdots\times X_{n}$, $L=p_{1}^{*}(L_{1})\otimes\cdots\otimes p_{n}^{*}(L_{n})$ and $f=f_{1}\times\cdots\times f_{n}$, that is, $f(x_{1},\ldots,x_{n})=(f_{1}(x_{1}),\ldots,f_{n}(x_{n}))$ for $(x_{1},\ldots,x_{n})\in X$. Moreover, let $p_{i}:X\to X_{i}$ be the projection to the $i$-th factor. Note that $f^{*}(L)\simeq p_{1}^{*}(L_{1})^{\otimes d_{1}}\otimes\cdots\otimes p_{n}^{*}(L_{n})^{\otimes d_{n}}$, and $\varphi=p_{1}^{*}(\varphi_{1})\otimes\cdots\otimes p_{n}^{*}(\varphi_{n})$ yields the global compatification of $f^{*}(L)\simeq p_{1}^{*}(L_{1})^{\otimes d_{1}}\otimes\cdots\otimes p_{n}^{*}(L_{n})^{\otimes d_{n}}$, that is, we have the isometry
$f_{\omega}^{*}(L_{\omega},\varphi_{\omega})\simeq p_{1,\omega}^{*}(L_{1,\omega},\varphi_{1,\omega})^{\otimes d_{1}}\otimes\cdots\otimes p_{n,\omega}^{*}(L_{n,\omega},\varphi_{n,\omega})^{\otimes d_{n}}$
for each $\omega$.
Proposition 4.1.
For $x_{1}\in X_{1}(K),\ldots,x_{n}\in X_{n}(K)$,
$\begin{cases}h_{f}(x_{1},\ldots,x_{n})=h_{f_{1}}(x_{1})+\cdots+h_{f_{n}}(x_{n}),\\ h_{f}(f(x_{1},\ldots,x_{n}))=d_{1}h_{f_{1}}(x_{1})+\cdots+d_{n}h_{f_{n}}(x_{n}).\end{cases}$
Proof.
Indeed,
$h_{f}(x_{1},\ldots,x_{n})$ $=h_{(p_{1}^{*}(L_{1})\otimes\cdots\otimes p_{n}^{*}(L_{n}),p_{1}^{*}(\varphi_{1})\otimes\cdots\otimes p_{n}^{*}(\varphi_{n}))}(x_{1},\ldots,x_{n})$
$=h_{(p_{1}^{*}(L_{1}),p_{1}^{*}(\varphi_{1}))}(x_{1},\ldots,x_{n})+\cdots+h_{(p_{n}^{*}(L_{n}),p_{n}^{*}(\varphi_{n}))}(x_{1},\ldots,x_{n})$
$=h_{f_{1}}(x_{1})+\cdots+h_{f_{n}}(x_{n}).$
Moreover,
$h_{f}(f(x_{1},\ldots,x_{n}))$ $=h_{f}(f_{1}(x_{1}),\ldots,f_{n}(x_{n}))=h_{f_{1}}(f_{1}(x_{1}))+\cdots+h_{f_{n}}(f_{n}(x_{n}))$
$=d_{1}h_{f_{1}}(x_{1})+\cdots+d_{n}h_{f_{n}}(x_{n}),$
as required. ∎
Lemma 4.2.
Let $S$ be a finite subset of $X(K)$. Let $g_{i}:X_{i}\to X_{i}$ be an endomorphism of $X_{i}$ such that $g_{i}$ are compatible with $L_{i}$, $d(g_{i})\in{\mathbb{Z}}_{\geqslant 2}$, and $f_{i}\circ g_{i}=g_{i}\circ f_{i}$ for each $i\in\{1,\ldots,n\}$. We set $g=g_{1}\times\cdots\times g_{n}$. If $\min\{d(g_{i})a(h_{f_{i}})\mid i\in\{1,\ldots,n\}\}>\max\{h_{f}(s)\mid s\in S\}$ and $g(x)\in S$ for $x\in X(K)$, then $h_{f}(x)=0$.
Proof.
We set $x=(x_{1},\ldots,x_{n})$ and $I=\{i\in\{1,\ldots,n\}\mid h_{f_{i}}(x_{i})>0\}$. If we assume $I\not$$\emptyset$, then
$\max\{h_{f}(x)\mid x\in S\}$ $\geqslant h_{f}(g(x))=\sum_{i=1}^{n}h_{f_{i}}(g_{i}(x_{i}))=\sum_{i=1}^{n}h_{g_{i}}(g_{i}(x_{i}))$
$=\sum_{i=1}^{n}d(g_{i})h_{g_{i}}(x_{i})=\sum_{i=1}^{n}d(g_{i})h_{f_{i}}(x_{i})$
$=\sum_{i\in I}d(g_{i})h_{f_{i}}(x_{i})\geqslant\sum_{i\in I}d(g_{i})a(h_{f_{i}})$
$\geqslant\min\{d(g_{i})a(h_{f_{i}})\mid i\in\{1,\ldots,n\}\},$
which is a contradiction. ∎
For $I=(i_{1},\ldots,i_{n})\in({\mathbb{Z}}_{\geqslant 1})^{n}$ and $1\leqslant k\leqslant n$, we denote $i_{k}$ by $I(k)$. We denote $\min\{I(k)\mid k\in\{1,\ldots,n\}\}$ by $\min I$. For $I,J\in({\mathbb{Z}}_{\geqslant 1})^{n}$, we define $I\leqslant J$, $I+J$ and $I\cdot J$ to be
$\begin{cases}I\leqslant J\ \Longleftrightarrow\ I(1)\leqslant J(1),\ldots,I(n)\leqslant J(n),\\ I+J=(I(1)+J(1),\ldots,I(n)+J(n)),\\ I\cdot J=(I(1)\cdot J(1),\ldots,I(n)\cdot J(n)).\end{cases}$
Moreover, we say $I$ is prime if $I(k)$ is a prime number for every $k\in\{1,\ldots,n\}$.
Let $F_{1}=\{f_{1,N}\}_{N=N_{1}}^{\infty},\ldots,F_{n}=\{f_{n,N}\}_{N=N_{n}}^{\infty}$ be systems of endomorphisms on $X_{1},\ldots,X_{n}$, respectively such that $f_{k,N}$ is compatible with $L_{k}$, that is, $f^{*}_{k,N}(L_{k})\simeq L^{\otimes d_{k,N}}$ ($d_{k,N}\in{\mathbb{Z}}_{\geqslant 2}$). We set $I_{0}=(N_{1},\ldots,N_{n})$. For $I\geqslant I_{0}$, the endomorphism $f_{1,I(1)}\times\cdots\times f_{n,I(n)}$ on $X$ is denoted by $f_{I}$, and the collection $\{f_{I}\}_{I\in({\mathbb{Z}}_{\geqslant 1})^{n}_{\geqslant N_{0}}}$ is denoted by $F$. Note that the following properties are easily verified:
  1. (1)
    For $I,I^{\prime}\in({\mathbb{Z}}_{\geqslant 1})^{n}_{\geqslant N_{0}}$, $f_{I}\circ f_{I^{\prime}}=f_{I^{\prime}}\circ f_{I}$.
  2. (2)
    $f_{I}^{*}(p_{1}^{*}(L_{1})\otimes\cdots\otimes p_{n}^{*}(L_{n}))=p_{1}^{*}(L_{1})^{\otimes d_{1,I(1)}}\otimes\cdots\otimes p_{n}^{*}(L_{n})^{\otimes d_{n,I(n)}}$.
Moreover, note that the height function $h_{f_{I}}$ does not depend on the choice of $I$, so it is denoted by $h_{F}$. Let $Y$ be an equidimensional subscheme of $X$, and $Y_{I}=f^{-1}_{I}(Y)$. We say $Y_{I}$ has Fermat’s property over $K$ if
$Y_{I}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}.$
Corollary 4.3.
If $F$ is multiplicative, that is, $f_{I}\circ f_{I^{\prime}}=f_{I^{\prime}}\circ f_{I}=f_{I\cdot I^{\prime}}$ for all $I,I^{\prime}\in({\mathbb{Z}}_{\geqslant 1})^{n}_{\geqslant N_{0}}$ and $Y_{I}(K)$ is finite, then there is $M_{0}$ such that $Y_{I\cdot M}$ has Fermat’s property for all $M\geqslant M_{0}$.
Proof.
This is a consequence of Lemma 4.2. ∎
Conjecture 4.4 (Multi-indexed version of generalized Fermat’s conjecture).
If there is $I_{1}\geqslant I_{0}$ such that $Y_{I}(K)$ is finite for any $I\geqslant I_{1}$, then does there exist $I_{2}\geqslant I_{0}$ such that $Y_{I}$ has Fermat’s property over $K$ for any $I\geqslant I_{2}$?
One has the following as an evidence of the above conjecture.
Theorem 4.5.
If $F$ is multiplicative and there is $I_{1}\geqslant I_{0}$ such that $Y_{I}(K)$ is finite for all $I\geqslant I_{1}$, then
$\small\lim_{m_{1}\to\infty,\cdots,m_{n}\to\infty}\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$I_{0}\leqslant I\leqslant(m_{1},\ldots,m_{n})$ and $Y_{I}$ has Fermat's property over $K$}\}}{m_{1}\cdots m_{n}}=1.$}
Proof.
This is a consequence of the above corollary and the following lemma. ∎
Lemma 4.6.
Let $\Sigma$ be a subset of $({\mathbb{Z}}_{\geqslant 1})^{n}$. We assume that there is a positive integer $p_{0}$ with the following properties: for any prime $P\in({\mathbb{Z}}_{\geqslant 1})^{n}$ with $\min P\geqslant p_{0}$ (i.e., $P(i)$ is a prime number and $P(i)\geqslant p_{0}$ for every $i$), there is $M^{\prime}\in({\mathbb{Z}}_{\geqslant 1})^{n}$ such that $P\cdot M\in\Sigma$ for any $M\geqslant M^{\prime}$. Then
$\lim_{m_{1}\to\infty,\ldots,m_{n}\to\infty}\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$I\in\Sigma$ and $I\leqslant(m_{1},\ldots,m_{n})$}\}}{m_{1}\cdots m_{n}}=1.$
Proof.
Since the Riemann zeta function has a pole at $1$, for a positive number $\varepsilon$ with $0<\varepsilon\leqslant 1$, let $p_{1},\ldots,p_{\ell}$ be prime numbers such that $p_{0}\leqslant p_{1}<\cdots<p_{\ell}$ and $(1-1/p_{1})\cdots(1-1/p_{\ell})\leqslant\varepsilon/(4n)$. For each $Q\in\{p_{1},\ldots,p_{\ell}\}^{n}$, by our assumption, there exists $M_{Q}\in({\mathbb{Z}}_{\geqslant 1})^{n}$ such that $Q\cdot M\in\Sigma$ for all $M\geqslant M_{Q}$.
Claim 4.6.1.
If we set $e=p_{1}\cdots p_{\ell}$, then we can find $M_{0}\in{\mathbb{Z}}_{\geqslant 1}^{n}$ such that
$\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$I\geqslant M_{0}$ and $\operatorname{GCD}(I(i),e)\not=1$ for all $i\in\{1,\ldots,n\}$}\}\subseteq\Sigma$
Proof.
We set $M_{0}=\sum_{Q\in\{p_{1},\ldots,p_{\ell}\}^{n}}QM_{Q}$. Let $I$ be an element of $({\mathbb{Z}}_{\geqslant 1})^{n}$ such that $I\geqslant M_{0}$ and $\operatorname{GCD}(I(i),e)\not=1$ for all $i\in\{1,\ldots,n\}$. Then, for each $i\in\{1,\ldots,n\}$, there is $q_{i}\in\{p_{1},\ldots,p_{\ell}\}$ such that $q_{i}\mid I(i)$. Thus, if we set $Q=(q_{1},\ldots,q_{n})$, then $I=Q\cdot M$ for some $M\in({\mathbb{Z}}_{\geqslant 1})^{n}$. Moreover,
$QM_{Q}\leqslant M_{0}\leqslant I=Q\cdot M,$
so $M\geqslant M_{Q}$, and hence $I\in\Sigma$. ∎
We set $M_{0}=(a_{1},\ldots,a_{n})$. Then we have the following:
Claim 4.6.2.
If $m_{1}\geqslant a_{1},\ldots,m_{n}\geqslant a_{n}$, then
$\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\setminus\Sigma\mid I\leqslant(m_{1},\ldots,m_{n})\}}{m_{1}\cdots m_{n}}\leqslant\sum_{i=1}^{n}\frac{a_{i}-1}{m_{i}}+(1-1/p_{1})\cdots(1-1/p_{\ell})\sum_{i=1}^{n}\frac{m_{i}+e}{m_{i}}$
Proof.
By the above claim,
$({\mathbb{Z}}_{\geqslant 1})^{n}\setminus\Sigma$ $\subseteq\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid I\not\geqslant M_{0}\}\cup\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$\exists i\in\{1,\ldots,n\}$ $\operatorname{GCD}(I(i),e)=1$}\}$
$\kern-20.00003pt\subseteq\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$\exists i\in\{1,\ldots,n\}$ $I(i)<a_{i}$}\}\cup\bigcup_{i=1}^{n}\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$\operatorname{GCD}(I(i),e)=1$}\}$
$\kern-20.00003pt\subseteq\bigcup_{i=1}^{n}\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$I(i)<a_{i}$}\}\cup\bigcup_{i=1}^{n}\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$\operatorname{GCD}(I(i),e)=1$}\}$
Thus, by [8, the proof of Lemma 5.16], if $m_{1}\geqslant a_{1},\ldots,m_{n}\geqslant a_{n}$, then
$\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\setminus\Sigma\mid I\leqslant(m_{1},\ldots,m_{n})\}}{m_{1}\cdots m_{n}}\leqslant\sum_{i=1}^{n}\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$I(i)<a_{i}$ and $I\leqslant(m_{1},\ldots,m_{n})$}\}}{m_{1}\cdots m_{n}}\\ +\sum_{i=1}^{n}\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$\operatorname{GCD}(I(i),e)=1$ and $I\leqslant(m_{1},\ldots,m_{n})$}\}}{m_{1}\cdots m_{n}}\\ \leqslant\sum_{i=1}^{n}\frac{a_{i}-1}{m_{i}}+\sum_{i=1}^{n}\frac{(1-1/p_{1})\cdots(1-1/p_{\ell})(m_{i}+e)}{m_{i}},$
as required. ∎
We choose positive integers $b_{1},\ldots,b_{n}$ such that $(a_{i}-1)/b_{i}\leqslant\varepsilon/(2n)$ and $e/b_{i}\leqslant\varepsilon$. Then, by the above claim, if $m_{i}\geqslant\max\{a_{i},b_{i}\}$ for all $i\in\{1,\ldots,n\}$, we have
$\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\setminus\Sigma\mid I\leqslant(m_{1},\ldots,m_{n})\}}{m_{1}\cdots m_{n}}\leqslant\varepsilon/2+(\varepsilon/4)(1+\varepsilon)\leqslant\varepsilon,$
because $0<\varepsilon\leqslant 1$, and hence the assertion follows. ∎